Usually, when you purchase a lot in an auction, you pay not only your winning bid price but also a buyer's premium. At one auction house, the buyer's premium is 21% of the winning bid price for purchases up to $60,000. For larger purchases, the buyer’s premium is 21% of the first $60,000 plus 7% of the purchase price above $60,000. Express the total purchase price of a lot at this auction house as a function of the final (winning) bid price. Sketch the graph of this function.
Answer #1

Let us denote the final (winning) bid price by x. Then, if x<=$60,000 the total amount to be paid will be equal to:

=x + 21% of x

=x (1+0.21)

= 1.21x

similarly if x>60,000 the total amount to be paid will be equal to:

= x + 21% of 60,000 + 7% of (x-60,000)

= x + 12600 + 0.07*(x-60000)

= x + 12600 + 0.07x - 4200

= 1.07x + 8400

So we can write the purchase price as a function of x as

p(x)= 1.21.2 , I < 60,000 11.07.0 + 8400 ,> 60,000

The graph will look like as shown below:

D+vC6OD8ynRmAAAAAElFTkSuQmCC